When the sum of 424 and three times a positive number is subtracted from the square of the number, the result is 126. Find the number.
Answer by ewatrrr(24785)

n^2 -(3n+424) = 126
n^2 - 3n - 550 = 0 (Tossing out the negative solution for unit measure)
(n-25)(n+22) = 0, n = 25